OpenStudy (anonymous):

Solve the equation. Check the solution. -2/x+4 = 4/x+3

OpenStudy (anonymous):

\[\frac{-2}{x+4}=\frac{4}{x+3}\iff 4(x+4)=-2(x+3)\] and go from there

OpenStudy (mathmale):

Satellite did you the courtesy of "cross-multiplying" that equation of fractions.

OpenStudy (anonymous):

I'm completely lost still >.<

OpenStudy (anonymous):

\[\frac{a}{b}=\frac{c}{d}\iff ad=bc\]

OpenStudy (anonymous):

Ohh

OpenStudy (anonymous):

solve the equation i wrote on the right hand side for \(x\)

OpenStudy (anonymous):

So it would be 4x +16 = -2x - 6 and then I would add 6 to both sides so 4x + 22 = -2x and then take 4x away from both sides 22 = -6x 22/-6 simplify 11/-3?

OpenStudy (mathmale):

Now, Dani: Why not check your result by substituting it back into the original equation? If the original eq'n is true when x=-11/3, then your solution is indeed a solution.